The line l joining the points M(3,4) and N(8,14) cuts the x-axis at A and the y-axis at B.
a. Find the general equation of the line l.
b. Find the coordinate of Point A.
c. Find the coordinate of Point B.

Answer by mananth(16949)

x1 y1 x2 y2
3 4 8 14

slope m =(y2-y1)/(x2-x1)
(14-4)/(8-3)
(10/5)
m=2

Plug value of the slope and point (3,4) in
Y = m x + b
4.00 = 6 + b
b= 4 - 6
b= -2
So the equation will be
Y = 2 x -2
-2x+y+2=0
This cuts the X axis at A so y=0
plug y=0
-2x=-2
/-2
x=1
A(1,0)
Plug x=0 for y intercept
y=-2
B=(0,-2)
m.ananth@hotmail.ca